# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v3

@@ Results
In Figure 2, we show exemplary results from standard NPE, our semi-supervised NPE ($M=8$), and Stan as reference. We see that standard NPE is highly inaccurate for many countries, whereas our semi-supervised approach is in strong agreement with the reference for all but one country. As shown in Table 1, adding the self-consistency loss ($M=8$) strongly improves posterior estimates for all five parameters across all metrics, on average across countries. The complete results can be found in Appendix A.5.
@@ 5 Discussion
In this paper, we demonstrated that Bayesian self-consistency losses can drastically improve the robustness of neural amortized Bayesian inference (ABI) on out-of-simulation data. Performing accurate inference on such data – outside of the space seen during simulation-based training (e.g., because of model misspecification) – has been one of the major challenges of ABI since its inception: conventional ABI approaches are known to dramatically fail in such cases [37, 19, 25], as we also illustrated in our experiments. In contrast, when adding the self-consistency loss and training it on unlabeled out-of-simulation data, we obtained near-unbiased posterior estimation far beyond the training data spaces. That is, ABI became highly accurate even for data sets far away from any data seen during training by either of the two loss components – the standard simulation-based loss and the self-consistency loss. Importantly, because the self-consistency loss does not require data labels (i.e., knowledge of true parameter values), we can also use any amount of real data during training to improve the robustness of ABI.
We believe that the self-consistency loss trained on out-of-simulation data is a major breakthrough in neural ABI. The self-consistency loss was initially proposed to improve training efficiency for models with slow simulators [38, 26], but it was not considered in combination with (potentially real) out-of-simulation data. To the best of our knowledge, no existing ABI approach achieves similar degrees of robustness as our method presented here. The strong robustness gains persisted even in relatively high dimensional models (tested up to 100 parameters), suggesting no major curse of dimensionality. Additionally, using a learned (i.e., approximate) rather than a known likelihood density also increased the robustness significantly. However, the resulting amortized inference with a learned likelihood did not (yet) achieve the same level of accuracy on out-of-simulation data as in the known likelihood case. Our variance-based self-consistency loss relies on fast density evaluations during training, keeping times competitive. This makes free-form methods such as flow matching [31] less practical due to their need for numerical integration. As a result, joint learning of posteriors and likelihoods, along with efficient self-consistency losses for free-form flows remains an open research challenge.
@@ Acknowledgments
Appendix A Appendix
A.1 Proofs
@@
A.2 Detailed setup of the multivariate normal case study
@@
A.3 Comprehensive results for the multivariate normal case study
Figure 3: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior for variations of the default configuration outlined in Section 4.1. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$).
Figure 4: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior when the likelihood is estimated (top row) and in presence of a summary network ($K=10$ data points; bottom row). In the setting where the likelihood function is estimated, we observe a lower bias of the posterior mean and lower maximum mean discrepancy to the true posterior when the self-consistency loss component is added compared to the standard simulation-based loss alone. However, we do see some bias of the posterior standard deviation, although with reversed signed compared to the standard loss. The self-consistency loss provides strong robustness gains in the presence of a summary network (and known likelihood) in terms of all metrics.
A.4 Detailed setup of the air traffic case study
@@
A.5 Comprehensive results for the air traffic case study
@@
$\alpha$ $0.078$ 0.003 $0.033$ 0.019 $0.085$ 0.032
$\beta$ $0.152$ 0.011 $0.054$ 0.011 $0.161$ 0.070
$\gamma$ $0.087$ 0.047 $0.057$ 0.022 $0.153$ 0.112
$\delta$ $0.052$ 0.046 $0.037$ 0.018 $0.119$ 0.102
$\log(\sigma)$ $0.214$ 0.207 0.049 $0.057$ $0.304$ 0.281
@@
$\alpha$ $0.078$ 0.002 $0.033$ 0.001 $0.085$ 0.006
$\beta$ $0.152$ 0.001 $0.054$ 0.001 $0.161$ 0.009
$\gamma$ $0.087$ 0.002 $0.057$ 0.006 $0.153$ 0.014
$\delta$ $0.052$ 0.003 $0.037$ 0.005 $0.119$ 0.013
$\log(\sigma)$ $0.214$ 0.002 $0.049$ 0.004 $0.304$ 0.011

### v4

@@ Results
In Figure 3, we show exemplary results from standard NPE, our semi-supervised NPE ($M=8$), and Stan as reference. We see that standard NPE is highly inaccurate for many countries, whereas our semi-supervised approach is in strong agreement with the reference for all but one country. As shown in Table 1, adding the self-consistency loss ($M=8$) strongly improves posterior estimates for all five parameters across all metrics, on average across countries. The complete results along with standard error values for Table 1 can be found in Appendix E.
4.3 Hodgkin-Huxley model of neuron activation
To investigate the effect of the self-consistency loss on a model involving high-dimensional data, we evaluate our approach on the Hodgkin-Huxley model, which was previously used in an ABI setting by Gloeckler et al. [20]. The Hodgkin-Huxley model is a classical model in neuroscience to describe neuron activation via a set of 5 ordinary differential equations. In brief, the model has 7 parameters (electrical conductances of different ion channels, membrane capacitance and reversal potentials), and the output $y_{i}$ with observation index $i$ is a 200-dimensional time series of the membrane potential. A full definition of the model, as well as the training setup and a description of the neural architectures, are shown in Appendix F.
To facilitate network training, all parameters are transformed to follow standard normal distributions through appropriate transformations. For example, the parameter $g_{\text{Na}}$ with marginal prior $g_{\text{Na}}\sim\text{LogNormal}(\log(110),0.1²)$ is transformed via $z_{g_{\text{Na}}}=(\log(g_{\text{Na}})-\log(110))/0.1$. We denote the full set of transformed model parameters by $\theta$. Training is performed with a simulation budget of $N=32{,}768$. For each loss evaluation, the self-consistency loss is computed on a random subset of 32 samples drawn from a pool of $M=1{,}024$ unlabeled observations. These are generated by first sampling $\theta\sim\text{Normal}(0,2)$, applying the inverse transformations to recover the original parameter scale, and then simulating time series of the membrane potential as above.
Results
(a) Posterior predictive samples without (top row) and with (bottom row) self-consistency loss.
(b) Quantitative evaluation of predictive bias.
Figure 4: (a) Posterior predictive samples (gray) inferred from an out-of-simulation dataset (black). NPE only produces highly biased predictions while NPE+SC yields accurate results. (b) Histogram of the mean absolute bias (MAB) difference of posterior predictions computed for 1000 out-of-simulation datasets. NPE+SC has lower bias than NPE for almost all datasets.
To assess the benefits of our approach in the out-of-distribution setting, Figure 11(a) shows posterior predictive samples inferred from data simulated with $\theta\sim\text{Normal}(-2,1)$. This contrasts with training data from $\theta\sim\text{Normal}(0,1)$ and self-consistency evaluation data from $\theta\sim\text{Normal}(0,2)$. The plot shows that, when training without the self-consistency loss, the neural posterior density estimator produces samples inconsistent with the observed data, while incorporating the loss yields accurate predictions (see Appendix G for further results). To quantify this, Figure 11(b) reports mean absolute bias differences between the two estimators. The self-consistency loss consistently and strongly improves predictions across the majority of time series. Even in the worst cases, it is at least competitive with the estimator trained without the self-consistency loss.
4.4 Bayesian denoising of MNIST images
Finally, we illustrate the utility of our self-consistency loss based semi-supervised approach in the high-dimensional setting of image denoising in a set-up similar to [10]. The parameter vector $\theta\in\mathbb{R}^{784}$ is the flattened image, and the observation $x\in\mathbb{R}^{784}$ is a blurry version of the same image generated by a simulated noisy camera. We assume an implicit prior $\theta\sim p(\theta)$ defined by a generative model trained on blurred MNIST images of digit “0” [31], and an implicit likelihood $p(x|\theta)$ implemented by reapplying the same amount of blur to each $\theta$. This creates a challenging neural posterior-likelihood estimation (NPLE) problem (see also Section 2.3).
To generate our training set, we first blur all digit “0” images in the MNIST training set with a fixed Gaussian filter and train a neural network to (i) sample new blurred simulated images $\{\theta^{i}\}_{i=1}^{N}\sim p(\theta)$ and (ii) evaluate their log-probabilities. For each sample $\theta^{i}$, we then produce an observation $x^{i}\sim p(x|\theta^{i})$ by reapplying the same amount of Gaussian blur. We generate $N=12000$ pairs $(\theta^{i},x^{i})$ to train our posterior and likelihood networks. For the self-consistency loss, we use a held-out subset of 400 MNIST test set images blurred only by the likelihood model (i.e., no additional prior blur involved) during the training. This deliberate mismatch induces a prior misspecification, allowing us to evaluate the robustness gains from using self-consistency loss in an NPLE setting. More details about the training set-up and model architecture can be found in the Appendix H.
Results
We perform inference on another held-out subset of MNIST test images comprising 580 images. Figure 5 (also Figure 14 in Appendix I) depicts ten randomly chosen examples alongside the posterior mean and standard deviation maps computed from 500 samples. Reconstructions using the self-consistency approach (NPLE+SC) are smoother and show more resemblance to the ground-truth image. In contrast, the means of posterior samples from NPLE are highly pixelated and blurry. Moreover, the standard deviation maps of NPLE+SC are far more coherent as elevated variations appear only along the inner and outer contours of the “0” precisely where one expects genuine edge ambiguity. In contrast, NPLE estimates exhibit scattered, patchy uncertainty across the digit and background, reflecting spurious standard deviation estimates. Figure 15 in Appendix I shows several individual posterior samples of seven randomly chosen images from the test-set for both NPLE+SC and NPLE approaches. The NPLE+SC posterior samples are smoother, less pixelated and better resemble the true image further reaffirming the advantage of including self-consistency loss.
Figure 5: Example of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts the same images after applying the Gaussian blur ($x$), third and fourth rows depict mean and SD of 500 posterior samples estimated from the corresponding blurry observations using NPLE+SC, and the fifth and sixth rows depict the mean and SD of 500 posterior samples using NPLE only. Incorporating SC loss significantly improves denoising: reconstructed means become smoother, less pixelated, and closer to the ground truth. In the standard-deviation maps, darker regions indicate higher output variability; NPLE + SC approach produces coherent maps with variability confined to the inner and outer edges.
@@ 5 Discussion
We demonstrated that Bayesian self-consistency losses significantly increase the robustness of neural amortized Bayesian inference (ABI) on out-of-simulation data. Accurate inference outside the training distribution, such as in the presence of model misspecification, has long posed a major challenge for ABI. While self-consistency was originally introduced to improve training efficiency with slow simulators [39, 26], it had not been previously explored as a remedy to simulation gaps. Existing supervised ABI approaches have been known to dramatically fail in such cases [38, 20, 25], as we also illustrated in our experiments. In contrast, when optimizing for self-consistency on unlabeled out-of-simulation data, we obtained nearly unbiased posterior estimation far beyond the training distribution. The strong robustness gains persisted even in models with several hundred parameters. Additionally, using a neural (i.e., approximate) in place of an analytic likelihood density also increased the robustness significantly. Finally, as self-consistency losses do not require data labels (i.e., true parameter values), we can use any amount of real data during training to improve the robustness of ABI.
Limitations and future directions
A notable limitation is that our variance-based self-consistency loss relies on fast density evaluations during training, keeping times competitive. This makes free-form methods such as flow matching [32] or score-based diffusion [41] less practical due to their need for numerical integration. As a result, efficient self-consistency losses for free-form flows, along with joint learning of posteriors and very high-dimensional likelihoods, remains an open avenue for future research.
@@ Acknowledgments
Appendix
Appendix A Proofs
@@
Appendix B Detailed setup of the multivariate normal case study
@@
Appendix C Comprehensive results for the multivariate normal case study
Figure 6: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior for variations of the default configuration outlined in Section 4.1. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$). Errorbars show ±$\pm 1$ standard deviations over 10 model refits on new training data.
Figure 7: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior when the likelihood is estimated (top row) and in presence of a summary network ($K=10$ data points; bottom row). In the setting where the likelihood function is estimated, we observe a lower bias of the posterior mean and lower maximum mean discrepancy to the true posterior when the self-consistency loss component is added compared to the standard simulation-based loss alone. However, we do see some bias of the posterior standard deviation, although with reversed signed compared to the standard loss. The self-consistency loss provides strong robustness gains in the presence of a summary network (and known likelihood) in terms of all metrics. Errorbars show ±$\pm 1$ standard deviations over 10 model refits on new training data.
Appendix D Detailed setup of the air traffic case study
@@
Appendix E Comprehensive results for the air traffic case study
@@
$\alpha$ $0.079$ 0.003 $0.033$ 0.020 $0.086$ 0.033
$\beta$ $0.153$ 0.012 $0.055$ 0.011 $0.161$ 0.070
$\gamma$ $0.087$ 0.048 $0.058$ 0.022 $0.154$ 0.112
$\delta$ $0.053$ 0.046 $0.038$ 0.018 $0.119$ 0.102
$\log(\sigma)$ $0.215$ 0.207 0.049 $0.058$ $0.304$ 0.282
@@
$\alpha$ $0.079$ 0.002 $0.033$ 0.002 $0.086$ 0.006
$\beta$ $0.153$ 0.001 $0.055$ 0.001 $0.161$ 0.009
$\gamma$ $0.087$ 0.002 $0.058$ 0.005 $0.154$ 0.014
$\delta$ $0.053$ 0.003 $0.038$ 0.005 $0.119$ 0.013
$\log(\sigma)$ $0.215$ 0.002 $0.049$ 0.004 $0.304$ 0.011
[…]
Parameter $\text{SE}(|\mu-\mu_{\rm Stan}|)$ $\text{SE}(|\sigma-\sigma_{\rm Stan}|)$
NPE SC(4) SC(8) SC(15) NPE SC(4) SC(8) SC(15)
$\alpha$ $0.019$ $0.009$ $0.013$ 0.001 $0.011$ $0.008$ $0.007$ 0.001
$\beta$ $0.024$ $0.026$ $0.023$ 0.002 $0.010$ $0.007$ $0.003$ 0.002
$\gamma$ $0.045$ $0.037$ $0.026$ 0.003 $0.015$ $0.016$ $0.013$ 0.003
$\delta$ $0.033$ $0.033$ $0.023$ 0.004 $0.015$ $0.016$ $0.012$ 0.002
$\log(\sigma)$ $0.076$ $0.076$ $0.058$ 0.002 $0.015$ $0.017$ $0.008$ 0.002
@@
Appendix F Detailed setup of the neuron activation case study
F.1 Model description
The prior and likelihood are given by
$\begin{gathered}g_{\text{Na}}\sim\text{LogNormal}(\log(110),0.1^{2}),\,g_{\text{K}}\sim\text{LogNormal}(\log(36),0.1^{2}),\,g_{\text{M}}\sim\text{LogNormal}(\log(0.2),0.5^{2})\\ E_{\text{Na}}\sim\text{Normal}(50,5^{2}),\,E_{\text{K}}\sim\text{Normal}(-77,5^{2}),\,E_{\text{leak}}\sim\text{Normal}(-55,5^{2}),\,C_{m}\sim\text{Normal}(1,0.05^{2})\\ y_{i,t}\sim\text{Student-t}(V_{m}(t),0.1^{2},\text{df=10}),\end{gathered}$
[…]
The membrane voltage $V_{m}(t)$ evolves according to the classical Hodgkin-Huxley model, extended with an additional slow (muscarinic, M-type) potassium channel with current $I_{M}$. The total current across the membrane is modeled as:
$C_{m}\frac{dV_{m}}{dt}=I_{\text{Na}}+I_{\text{K}}+I_{\text{M}}+I_{\text{leak}}+I_{\text{in}}(t),$ (21)
where $I_{\text{Na}},I_{\text{K}}$, $I_{\text{M}}$, $I_{\text{leak}}$ are the sodium, potassium, M-type potassium, and leak currents respectively; and $I_{\text{in}}(t)$ is an externally applied current, which is set to be a pulse input of 3.248 nA between $t_{\text{on}}=10\text{ms}$ and $t_{\text{off}}=50\text{ms}$. The ionic currents are calculated as:
$I_{\text{Na}}$ $=g_{\text{Na}}m^{3}h(E_{\text{Na}}-V_{m})$ (22)
$I_{\text{K}}$ $=g_{\text{K}}n^{4}(E_{\text{K}}-V_{m})$ (23)
$I_{\text{M}}$ $=g_{\text{M}}p(E_{\text{K}}-V_{m})$ (24)
$I_{\text{leak}}$ $=g_{\text{leak}}(E_{\text{leak}}-V_{m}),$ (25)
where the leak conductance is fixed to $g_{\text{leak}}=0.1\text{mS/cm}^{2}$; and $m$, $h$, $n$, $p$ are gating variables. The gating variables take the form:
$\frac{dx}{dt}=\frac{x_{\infty}(V_{m})-x}{\tau_{x}(V_{m})},\qquad x\in\{n,m,h,p\},$ (26)
with $x_{\infty}(V_{m})=\alpha_{x}(V_{m})/(\alpha_{x}(V_{m})+\beta_{x}(V_{m}))$, and $\tau_{x}(V_{m})=1/(\alpha_{x}(V_{m})+\beta_{x}(V_{m}))$ for $n$, $m$ and $h$, where $\alpha_{x}$ and $\beta_{x}$ are voltage-dependent rate functions defined as:
$\alpha_{n}(V_{m})$ $=\frac{0.032\cdot\exp(-0.2(V_{m}-75))}{0.2},\quad$ $\beta_{n}(V_{m})$ $=\frac{0.28\cdot\exp(0.2(V_{m}-100))}{0.2},$
$\alpha_{m}(V_{m})$ $=\frac{0.32\cdot\exp(-0.25(V_{m}-73))}{0.25},\quad$ $\beta_{m}(V_{m})$ $=\frac{0.28\cdot\exp(0.2(V_{m}-100))}{0.2},$
$\alpha_{h}(V_{m})$ $=0.128\cdot\exp(\frac{-(V_{m}-77)}{18}),\quad$ $\beta_{h}(V_{m})$ $=\frac{4}{1+\exp(-0.2(V_{m}-100))}$
For the gating variable $p$ of the M-type potassium channel, a sigmoidal steady-state activation and custom time constant are used:
$p_{\infty}(V_{m})$ $=\frac{1}{1+\exp(-0.1(V_{m}+35))},\quad$ $\tau_{p}(V_{m})$ $=\frac{600}{3.3\cdot\exp(0.05(V_{m}+35))+\exp(-0.05(V_{m}+35))}$
[…]
F.2 Network architecture and training
For the NPEs $q(\theta|y_{i,t})$, we use a neural spline flow [8] with 10 coupling layers of 256 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-}3$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the summary network, we use a long short-term memory layer with 100 output dimensions followed by a sequence of dense layers with output dimensions of 400, 200, 100, and 50, respectively. The inference and summary network are jointly trained using the Adam optimizer with a batch size of 256 for 100 epochs and a fixed learning rate of $5\times 10^{-4}$, followed by 100 epochs with a fixed learning rate of $5\times 10^{-5}$ and a final run of 100 epochs with a learning rate of $5\times 10^{-6}$.
Appendix G Comprehensive results of the neuron activation case study
(a) Posterior predictive samples without (top row) and with (bottom row) self-consistency loss.
(b) Quantitative evaluation of predictive bias.
Figure 11: (a) Posterior predictive samples (gray) inferred from an in-simulation dataset (black) with parameters $\theta\sim\text{Normal}(0,1)$. Both NPE only and NPE+SC produce predictions that are consistent with the observed data. However, samples from NPE+SC are much closer to the ground truth. (b) Histogram of the mean absolute bias (MAB) difference of posterior predictions computed for 1000 out-of-simulation datasets. NPE+SC has lower bias than NPE for almost all datasets. Mean absolute bias is defined as $\text{MAB}(y_{i,t},\hat{y}_{i,t})=\frac{1}{T}\sum_{t=1}^{T}|y_{i,t}-\hat{y}_{i,t}|$ for a time series $y_{i,t}$ with observation index $i$ at time $t=1,\dots,T$. $\hat{y}_{i,t}=\frac{1}{S}\sum_{s=1}^{S}p(y_{i,t}|\theta^{(s)})$ denotes the mean of the posterior predictive distribution at time $t$ computed over $S$ posterior samples.
Figure 12: Posterior predictive samples (gray) inferred from 5 simulation datasets following the same distribution as the training data. Both NPE only and NPE+SC show predictions that are consistent with the observed data.
Figure 13: Posterior predictive samples (gray) inferred from 5 out-of-simulation datasets, generated from parameter draws $\theta\sim\text{Normal}(-2,1)$. While NPE only produces highly biased predictions, NPE+SC is consistent with the observed data.
Appendix H Detailed setup of the MNIST image denoising case study
We implement the jointly amortized posterior and likelihood networks [37] using two normalizing flows with fully connected affine-coupling layers that operate on the flattened 784-pixel vectors. We use the same network architecture that was used by [37]. As both $\theta$ and $x$ are images whose intrinsic dimensionality is significantly lower than their raw pixel count, we use identical 4-layer convolutional neural networks as summary networks for both posterior and likelihood networks. These summary networks terminate in a global average-pooling layer to produce a 128-dimensional summary of the original or blurred image, respectively. The posterior network itself is implemented as a conditional invertible neural network (cINN) consisting of 12 conditional affine-coupling layers; each coupling layer embeds its conditional information via an internal fully connected network with a single hidden layer of 512 units and ReLU activations. The likelihood network adopts exactly the same conditional-coupling architecture. In both cases, we employ a multivariate Student-T distribution in the latent space [1, 37], which enables more stable maximum-likelihood training at elevated learning rates.
The prior network that was used to generate blurred MNIST simulations followed the same architecture as the posterior network defined above. We applied the Gaussian blur with $\text{PSF}=1.0$ to the $5,923$ images of digit “0” in the MNIST training dataset to train the prior network using Adam optimizer for 120 epochs with a batch size of 32, learning rate of $1\times 10^{-3}$, and a $15\%$ dropout. After training, we generated $12000$ blurred images ($\theta$) of the digit 0 to train the posterior and likelihood networks. A Gaussian blur with $\text{PSF}=1.0$ was further applied to these images to generate observations ($x$) which represent images from a noisy camera. The posterior and likelihood networks along with summary networks were jointly trained on $\{\theta^{i},x^{i}\}_{i=1}^{12000}$ pairs for $100$ epochs with a batch size of $32$ using a learning rate of $1\times 10^{-4}$, and a $15\%$ dropout.
For self-consistency loss, the MNIST test set of digit “0” was divided into two subsets comprising 400 and 580 images respectively. The subset with 400 images was used for training self-consistency loss. A Gaussian blur with $\text{PSF}=1.0$ was applied to these images to generate observations ($x^{*}$). No prior blur was applied to these images. This represents a prior misspecification scenario as the simulated images used to train NPLE were already blurred before applying the noisy camera while the MNIST images used for inference do not have a prior blur. This misspecification scenario depicts the effectiveness of utilising self-consistency loss to overcome prior misspecification. The self-consistency loss was activated at epoch 21, with its weight linearly ramped from zero to one by epoch 40. Training was performed using minibatches of 16 images, and 32 consistency samples were drawn to estimate the variance. The inference was performed on the other held-out subset comprising 580 images and all the results in Figures 5, 14 and 15 use the MNIST images from this subset.
Appendix I Comprehensive results of the MNIST image denoising case study
[…]
[0.5ex]0.4pt3mm Refer to caption Refer to caption
Figure 14: More examples of denoising results for MNIST images of digit “0” in the held-out test set. The first row shows ten randomly selected MNIST images ($\theta$), the second row depicts the same images after applying the Gaussian blur ($x$), third and fourth rows depict the mean and standard deviation of 500 posterior samples estimated from the corresponding blurry observations using NPLE + SC based model, and the fifth and sixth rows depict the mean and standard deviation of 500 posterior samples from model based on NPLE only. Incorporating self-consistency loss significantly improves denoising as the means of reconstructed unblurred image are smoother, less-pixelated and better resemble the ground truth. The darker regions in the standard deviation show the regions of higher variability in the outputs. The standard deviation maps of NPLE + SC based approach are far more coherent showing high variability only along the inner and outer edges.
Figure 15: Ground-truth images and the corresponding posterior draws for seven randomly selected MNIST “0” digits from the held-out MNIST test set. In each row, the leftmost panel shows the true image ($\theta$), and the following panels show ten independent samples from the approximate posterior. The top-figure shows the posterior draws using the standard NPLE based model and the bottom figure shows posterior draws from combining self-consistency loss to the NPLE based model. It can clearly be seen that NPLE+SC posterior draws are a better reconstruction of the original image whereas NPLE based posterior samples are highly pixelated.
Appendix J Computational resources for experiments
  1. 1.
    Multivariate normal model: The experiments were run on a 16-core AMD Ryzen 5950x CPU, equipped with 32 GB of system RAM.
  2. 2.
    Air traffic case study: The experiments were conducted on a single MacBook Pro (M3, 2024) equipped with Apple’s M3 chip and 16 GB of unified RAM, running macOS Sonoma 14.6. We did not use the GPU cores.
  3. 3.
    Neuron activation case study: The experiments were run on a 16-core AMD Ryzen 5950x CPU, equipped with 32 GB of system RAM.
  4. 4.
    MNIST image denoising: The experiments were run on a high-performance compute cluster using a GPU-equipped compute node featuring a single NVIDIA Tesla P100-PCIE with 12 GB of dedicated HBM2 memory, paired with 16 GB of system RAM. The training for the longest experiment took $\sim 100$ minutes.